OpenStudy (anonymous):

24+6-2=12 add one thing to left side of the equation to have it work. P.S you cant give a negative number

OpenStudy (anonymous):

you want to get them equal?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2*4+6-2=12

OpenStudy (anonymous):

not split something up

OpenStudy (anonymous):

alright how about : 24 +6 -2 = 12 +16?

OpenStudy (anonymous):

to the left side

OpenStudy (anonymous):

not the right

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

24+6 - 2 -16=12

OpenStudy (anonymous):

no negatives

OpenStudy (anonymous):

you said hint: you can add a negative number

OpenStudy (anonymous):

lol, okay I got it ^^"

OpenStudy (anonymous):

ok what

OpenStudy (anonymous):

24 + 6 -(16) -2 ^_^ I didn't add a negative number I replaced it with 2 !

OpenStudy (anonymous):

or: 24 + 6 -(18) = 12 !

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

sqrt(24*6) -2+2 = 12 sqrt (144)=12

OpenStudy (anonymous):

is it right?